OpenStudy (anonymous):

Find the slope of the line passing through the points (-3,4) and (2,-1) 1. -5/3 2. 1 3. 3/5 4. -1

OpenStudy (alexandervonhumboldt2):

slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4-(-1) }{ -3-2 }=?\]

OpenStudy (anonymous):

3/5?

OpenStudy (solomonzelman):

4- (-1) = ? -3 -2 = ?

OpenStudy (solomonzelman):

You probably know that if you have, 7-(-6), then two minuses become a plus, and you would have 7+6, which is =13

OpenStudy (solomonzelman):

Same way, when you have, 4 - (-1), you get?

OpenStudy (anonymous):

so -4- (-1) is -5

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

first, -4 - (-1) = -4 + 1 =-3 And second, you need not that! you need; 4 - (-1) (without minus before 4)

OpenStudy (solomonzelman):

Yes, that is right it is 5 4 - (-1) = 5

OpenStudy (anonymous):

okay so its -3/5

pooja195 (pooja195):

\[\frac{ 4-(-1) }{ -3-2 }=?\] -3/5 is incorrect 2 negatives next to each other make a positive

pooja195 (pooja195):

so 4+1= ?

OpenStudy (anonymous):

5